# Confluent functions, Laguerre polynomials and their (generalized) bilinear integrals

### v1

@@ 3.1 Basic properties
$L_{n}^{\alpha}(z)$ $=\frac{(1+\alpha)_{n}}{n!}F(-n;1+\alpha;z)$ (3.3)
$=\frac{\Gamma(1+\alpha+n)}{n!}{\bf F}_{-1-\alpha-2n,\alpha}(z)$ (3.4)
$=\frac{(-z)^{n}}{n!}{}_{2}F_{0}(-n,-\alpha-n;-;-z^{-1})$ (3.5)
$=\frac{(-1)^{n}}{n!}U_{-1-\alpha-2n,\alpha}(z).$ (3.6)
@@
${\rm e}^{-tz}(1+t)^{\alpha}=\sum\limits_{n=0}^{\infty}t^{n}L_{n}^{\alpha-n}(z),$ (3.7)
$(1-t)^{-\alpha-1}\exp{\frac{tz}{t-1}}=\sum\limits_{n=0}^{\infty}t^{n}L_{n}^{\alpha}(z).$ (3.8)
@@
$L_{n}^{\alpha}(z)$ $=\frac{1}{2\pi{\rm i}}\int\limits_{[0^{+}]}{\rm e}^{-tz}(1+t)^{\alpha+n}t^{-n-1}{\rm d}t.$ (3.9)
@@
$L_{n}^{\alpha}(0)=\frac{(\alpha+1)_{n}}{n!},\ \ \ \ \lim\limits_{z\to\infty}\frac{L_{n}^{\alpha}(z)}{z^{n}}=\frac{(-1)^{n}}{n!}.$ (3.10)
@@
$L_{n}^{-\alpha}(z)$ $=\frac{(n-\alpha)!}{n!}(-z)^{\alpha}L_{n-\alpha}^{\alpha}(z),\quad\alpha\in{\mathbb{N}},\;\alpha\leq n.$ (3.11)
@@ 3.2 Bilinear integrals
Theorem 3.1.
Let $\mathrm{Re}(\alpha)>-1$. Then
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n}.$ (3.12)
@@
For completeness we sketch the proof. Assume without loss of generality that $m\geq n$. Applying the Rodriguez formula (3.1) for $L_{m}^{\alpha}$, and then integrating $m$ times by parts and using $\mathrm{Re}(\alpha)>-1$, one obtains
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\int_{0}^{\infty}\Big{(}\partial_{z}{\rm e}^{-z}z^{m+\alpha}\Big{)}L_{n}^{\alpha}(z){\mathrm{d}}z$ (3.13)
$=\frac{(-1)^{m}}{m!}\int_{0}^{\infty}{\rm e}^{-z}z^{m+\alpha}\partial_{z}^{m}L_{n}^{\alpha}(z){\mathrm{d}}z.$ (3.14)
Because $m\geq n$, we find $\partial_{z}^{m}L_{n}^{\alpha}(z)=(-1)^{m}\delta_{m,n}$ with the Kronecker delta on the right-hand side. Then the remaining integral is nothing but the integral representation of $\Gamma(1+n+\alpha)$. $\Box$
If we replace the integral on the left-hand side of (3.12) with the generalized integral, formula (3.12) remains valid if $\alpha\in{\mathbb{C}}\setminus-{\mathbb{N}}$, because then the generalized integral is non-anomalous and can be computed by analytic continuation:
Theorem 3.2.
Let $\alpha\in{\mathbb{C}}\setminus-{\mathbb{N}}$. Then (3.12) remains true if we replace the usual integral with the generalized integral.
@@
Theorem 3.3.
Let $\alpha\in-{\mathbb{N}}$ and $m,n\geq|\alpha|$. Then (3.12) remains true (in the sense of usual integrals):
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n}=\frac{(n-|\alpha|)!}{n!}\delta_{m,n}$ (3.15)
@@
By the identity (3.11)
$\quad\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (3.16)
@@
Theorem 3.4.
Let $-\alpha\in{\mathbb{N}}$, $m,n\in{\mathbb{N}}_{0}$ such that $|\alpha|>n$ and, without loss of generality, $m\geq n$. Then
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{(-1)^{\alpha+n}}{n!(|\alpha|-n-1)!(m-n)},\qquad n<m;$ (3.17)
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)^{2}z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{(-1)^{\alpha+n+1}}{n!(|\alpha|-n-1)!}\psi(|\alpha|-n).$ (3.18)
@@
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{1}{m!}\mathrm{gen}\int_{0}^{\infty}\big{(}\partial^{m}{\rm e}^{-z}z^{m+\alpha}\big{)}L_{n}^{\alpha}(z){\mathrm{d}}z.$ (3.19)
We may integrate the right-hand side by parts $n+1$ times, each time collecting the regular value of the remaining integrand. Thus
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{1}{m!}\sum_{k=1}^{n+1}(-1)^{k}\mathop{\mathrm{rv}_{0}}\Big{(}\big{(}\partial^{m-k}{\rm e}^{-z}z^{m+\alpha}\big{)}\partial^{k-1}L_{n}^{\alpha}(z)\Big{)}.$ (3.20)
@@
$\partial_{z}^{m-k}{\rm e}^{-z}z^{m+\alpha}$ $=(-1)^{m-k}\Bigg{(}\sum_{j=0}^{|\alpha|-m-1}+\sum_{j=|\alpha|-k}^{\infty}\Bigg{)}\frac{(-1)^{j}}{j!}(|\alpha|-m-j)_{m-k}z^{j+k-|\alpha|},$ (3.21)
$\partial_{z}^{k-1}L_{n}^{\alpha}(z)$ $=(-1)^{n}\sum_{i=0}^{n+1-k}\frac{(-\alpha-n)_{n-i-k+1}z^{i}}{i!(n-i-k+1)!},$ (3.22)
@@
$\quad\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (3.23)
@@
×$\quad\times\frac{(|\alpha|-n)_{n-k-i+1}}{i!(n-i-k+1)!}z^{i+k-|\alpha|+j}\Bigg{)}.$
The regular value evaluated at 00 is then the coefficient for $j=|\alpha|-k-i$. One quickly finds that only the term $i=0$ is non-zero, so actually, the regular value is the coefficient corresponding to $j=|\alpha|-k$ and $i=0$. Hence we are left with the single sum
[…]
@@
where in the last step we used an identity about telescoping series recalled in Lemma B.1.
Now consider $m=n<|\alpha|.$ We again use the Rodriguez type formula. However, we can now only integrate by parts $n$ times, so that a simple generalized integral survives:
$\quad\mathrm{gen}\int_{0}^{\infty}\big{(}L_{n}^{\alpha}(z)\big{)}^{2}z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (3.25)
@@
$\mathrm{gen}\int_{0}^{\infty}{\rm e}^{-z}z^{n+\alpha}{\mathrm{d}}z=-\frac{(-1)^{\alpha+n}}{(|\alpha|-n-1)!}\psi(|\alpha|-n),\quad n<|\alpha|\in{\mathbb{N}}.$ (3.26)
@@
$\quad\frac{1}{n!}\sum_{k=1}^{n}\mathop{\mathrm{rv}_{0}}\Bigg{(}\Bigg{(}\sum_{j=0}^{|\alpha|-n-1}+\sum_{j=|\alpha|-k}^{\infty}\Bigg{)}\sum_{i=0}^{n+1-k}\frac{(-1)^{j}}{j!}(|\alpha|-n-j)_{n-k}$ (3.27)
×$\quad\times\frac{(|\alpha|-n)_{1+n-k-i}}{i!(n-i-k+1)!}z^{i+k-|\alpha|+j}\Bigg{)}.$
@@
$G(\alpha)$ $=\operatorname{diag}\Big{(}\frac{\Gamma(1+\alpha+n)}{n!}\Big{)},$ $\quad\alpha\in{\mathbb{C}}\setminus-{\mathbb{N}},$ (3.29)
@@
In particular, all entries of the rows and columns $10,\dots,|\alpha|-$ 1 are non-zero and the sign of the entries of these rows and columns oscillates.
Integration by parts seems to be the simplest way to determine bilinear generalized integrals of Laguerre polynomials. Alternatively, we can derive Thms 3.2, 3.3 and 3.4 from Thm 2.2, using the fact that Laguerre polynomials are essentially special cases of Tricomi functions. For Thms 3.2 and 3.3 this is easy. To derive Thm 3.4 we need to analyze certain seemingly singular expressions.
Alternative proof of Thm. 3.4.
@@
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)L_{m}^{\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (3.30a)
$=$ $\frac{(-1)^{m+n}}{m!n!}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n,\alpha}(z)U_{-1-\alpha-2m,\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z,\quad m\neq n,$ (3.30b)
@@
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ (3.31a)
$=\frac{1}{(n!)^{2}}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z.$ (3.31b)
The generalized integrals of Tricomi functions given by Thm 2.2 are seemingly singular for the parameters in (3.30b) and (3.31b). However, these singularities are merely apparent. To see this, we replace (3.30b) and (3.31b) by the following limits:
$I_{1}$ $\coloneqq\frac{(-1)^{m+n}}{m!n!}\lim_{\epsilon\to 0}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n-2\epsilon,\alpha}(z)U_{-1-\alpha-2m-2\epsilon,\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z,\quad m\neq n,$ (3.32)
$I_{2}$ $\coloneqq\frac{1}{(n!)^{2}}\lim_{\epsilon\to 0}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n-2\epsilon,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z.$ (3.33)
@@
×$\times\Big{(}-\psi(|\alpha|-k)-\psi(k+1)+\tfrac{1}{2}H_{k}\big{(}\tfrac{1+\theta_{1}-|\alpha|}{2}\big{)}-\tfrac{1}{2}H_{|\alpha|-1-k}\big{(}\tfrac{3+\theta_{2}-|\alpha|}{2}+k\big{)}\Big{)}\Bigg{|}<\infty$ (3.34)
for any values of the parameters because possible vanishing denominators of the harmonic numbers are balanced by the Pochhammer symbol. Due to the $\Gamma$ functions in the prefactor, see Thm. 2.2, this sum only contributes if $-\alpha>m$ and $-\alpha>n$. In the cases where the sum does not contribute, we have
$I_{1}$ $=\frac{(-1)^{\alpha+m+n}}{2(m-n)n!m!}$ (3.35)
×$\quad\times\lim_{\epsilon\to 0}\begin{cases}\dfrac{\psi(-n-\epsilon)+\psi(-\alpha-n-\epsilon)}{\Gamma(-\alpha-n-\epsilon)\Gamma(-m-\epsilon)}-(m\leftrightarrow n),&\quad\alpha\geq 0;\vspace{12pt}\\ \dfrac{\psi(-n-\epsilon)+\psi(|\alpha|-n-\epsilon)}{\Gamma(-n-\epsilon)\Gamma(|\alpha|-m-\epsilon)}-(m\leftrightarrow n),&\quad\alpha<0,\;m\geq|\alpha|.\end{cases}$
@@
$I_{1}$ =$=\begin{cases}0,&\quad\alpha\geq 0;\vspace{12pt}\\ \dfrac{(-1)^{\alpha+n}}{(m-n)n!\Gamma(|\alpha|-n)},&\quad\alpha<0,\;m\geq|\alpha|.\end{cases}$ (3.37)
@@
Let us now assume that $-\alpha>m$ and $-\alpha>n$ (and still, $m>n$). Then
$I_{1}$ $=\frac{(-1)^{\alpha+m+n}}{2n!m!\Gamma(|\alpha|-m)\Gamma(|\alpha|-n)}\Bigg{(}\frac{(-1)^{m}m!\Gamma(|\alpha|-m)-(-1)^{n}n!\Gamma(|\alpha|-n)}{m-n}$ (3.38)
@@
$(z)_{k}H_{k}(z)=\sum_{j=0}^{k-1}\prod_{l=0,l\neq j}^{k-1}(z+l).$ (3.39)
With similar manipulations as in the proof of Theorem 3.4 and using Lemma B.1, the sum in the last line of (3.38) can be simplified and we obtain the expression from Thm 3.4.
The analysis of (3.33) is similar. We distinguish the three cases $\alpha>0$, $(\alpha<0\land n\geq|\alpha|)$ and $-\alpha>n$. The sum in (2.45) only contributes if $-\alpha>n$. To obtain the limit, one needs to evaluate certain combinations of $\psi^{\prime}(z)$, $\psi(z)$ and $\Gamma(z)$ at negative integers. We derive the respective formulas in Lemma B.2. $\Box$

### v2

@@ 3.1. Basic properties
$L_{n}^{\alpha}(z)$ $=\frac{(1+\alpha)_{n}}{n!}F(-n;1+\alpha;z)$ (72)
$=\frac{\Gamma(1+\alpha+n)}{n!}{\bf F}_{-1-\alpha-2n,\alpha}(z)$ (73)
$=\frac{(-z)^{n}}{n!}{}_{2}F_{0}(-n,-\alpha-n;-;-z^{-1})$ (74)
$=\frac{(-1)^{n}}{n!}U_{-1-\alpha-2n,\alpha}(z).$ (75)
@@
${\rm e}^{-tz}(1+t)^{\alpha}=\sum\limits_{n=0}^{\infty}t^{n}L_{n}^{\alpha-n}(z),$ (76)
$(1-t)^{-\alpha-1}\exp{\frac{tz}{t-1}}=\sum\limits_{n=0}^{\infty}t^{n}L_{n}^{\alpha}(z).$ (77)
@@
$L_{n}^{\alpha}(z)$ $=\frac{1}{2\pi{\rm i}}\int\limits_{[0^{+}]}{\rm e}^{-tz}(1+t)^{\alpha+n}t^{-n-1}{\rm d}t.$ (78)
@@
$L_{n}^{\alpha}(0)=\frac{(\alpha+1)_{n}}{n!},\ \ \ \ \lim\limits_{z\to\infty}\frac{L_{n}^{\alpha}(z)}{z^{n}}=\frac{(-1)^{n}}{n!}.$ (79)
@@
$L_{n}^{-\alpha}(z)$ $=\frac{(n-\alpha)!}{n!}(-z)^{\alpha}L_{n-\alpha}^{\alpha}(z),\quad\alpha\in{\mathbb{N}},\;\alpha\leq n.$ (80)
@@ 3.2. Bilinear integrals
Let $\Re(\alpha)>-1$. Then
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n}.$ (81)
@@
For completeness we sketch the proof. Assume without loss of generality that $m\geq n$. Applying the Rodrigues formula (70) for $L_{m}^{\alpha}$, and then integrating $m$ times by parts and using $\Re(\alpha)>-1$, one obtains
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\int_{0}^{\infty}\Big(\partial_{z}{\rm e}^{-z}z^{m+\alpha}\Big)L_{n}^{\alpha}(z){\mathrm{d}}z$ (82)
$=\frac{(-1)^{m}}{m!}\int_{0}^{\infty}{\rm e}^{-z}z^{m+\alpha}\partial_{z}^{m}L_{n}^{\alpha}(z){\mathrm{d}}z.$ (83)
Because $m\geq n$, we find $\partial_{z}^{m}L_{n}^{\alpha}(z)=(-1)^{m}\delta_{m,n}$ with the Kronecker delta on the right-hand side. Then the remaining integral is nothing but the integral representation of $\Gamma(1+n+\alpha)$. ∎
If we replace the integral on the left-hand side of (81) with the generalized integral, formula (81) remains valid if $\alpha\in{\mathbb{C}}\setminus$-$\mathbb{N}$}, because then the generalized integral is non-anomalous and can be computed by analytic continuation:
Let $\alpha\in{\mathbb{C}}\setminus$-$\mathbb{N}$}. Then (81) remains true if we replace the usual integral with the generalized integral.
@@
Let $\alpha\in$-$\mathbb{N}$} and $m,n\geq|\alpha|$. Then (81) remains true (in the sense of usual integrals):
$\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{\Gamma(1+n+\alpha)}{n!}\delta_{m,n}=\frac{(n-|\alpha|)!}{n!}\delta_{m,n}$ (84)
@@
By the identity (80)
$\quad\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (85)
@@
Let $\alpha\in{\mathbb{N}}$, $m,n\in{\mathbb{N}}_{0}$ such that $|\alpha|>n$ and, without loss of generality, $m\geq n$. Then
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{(-1)^{\alpha+n}}{n!(|\alpha|-n-1)!(m-n)},\qquad n<m;$ (86)
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)^{2}z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{(-1)^{\alpha+n+1}}{n!(|\alpha|-n-1)!}\psi(|\alpha|-n).$ (87)
@@
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ $=\frac{1}{m!}\mathrm{gen}\int_{0}^{\infty}\big(\partial^{m}{\rm e}^{-z}z^{m+\alpha}\big)L_{n}^{\alpha}(z){\mathrm{d}}z.$ (88)
We may integrate the right-hand side by parts $n+1$ times, each time collecting the regular value of the remaining integrand (cf. Appendix A for details). Thus
$\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z=\frac{1}{m!}\sum_{k=1}^{n+1}(-1)^{k}\mathop{\mathrm{rv}_{0}}\Big(\big(\partial^{m-k}{\rm e}^{-z}z^{m+\alpha}\big)\partial^{k-1}L_{n}^{\alpha}(z)\Big).$ (89)
@@
$\partial_{z}^{m-k}{\rm e}^{-z}z^{m+\alpha}$ $=(-1)^{m-k}\Bigg(\sum_{j=0}^{|\alpha|-m-1}+\sum_{j=|\alpha|-k}^{\infty}\Bigg)\frac{(-1)^{j}}{j!}(|\alpha|-m-j)_{m-k}z^{j+k-|\alpha|},$ (90)
$\partial_{z}^{k-1}L_{n}^{\alpha}(z)$ $=(-1)^{n}\sum_{i=0}^{n+1-k}\frac{(-\alpha-n)_{n-i-k+1}z^{i}}{i!(n-i-k+1)!},$ (91)
@@
$\quad\mathrm{gen}\int_{0}^{\infty}L_{m}^{\alpha}(z)L_{n}^{\alpha}(z)z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (92)
@@
$\quad\times\frac{(|\alpha|-n)_{n-k-i+1}}{i!(n-i-k+1)!}z^{i+k-|\alpha|+j}\Bigg).$
The regular value evaluated at 0 is then the coefficient for $j=|\alpha|-k-i$. One quickly finds that only the term $i=0$ is non-zero, so actually, the regular value is the coefficient corresponding to $j=|\alpha|-k$ and $i=0$. Hence we are left with the single sum
(92) $=\frac{(-1)^{m+n+|\alpha|-k}}{m!}\sum_{k=1}^{n+1}\frac{(k-m)_{m-k}(|\alpha|-n)_{n-k+1}}{(|\alpha|-k)!(n+1-k)!}$ (93)
@@
where in the last step we used an identity about telescoping series recalled in Lemma B.
Now consider $m=n<|\alpha|.$ We again use the Rodrigues type formula. However, we can now only integrate by parts $n$ times, so that a simple generalized integral survives:
$\quad\mathrm{gen}\int_{0}^{\infty}\big(L_{n}^{\alpha}(z)\big)^{2}z^{\alpha}{\rm e}^{-z}{\mathrm{d}}z$ (94)
@@
$\mathrm{gen}\int_{0}^{\infty}{\rm e}^{-z}z^{n+\alpha}{\mathrm{d}}z=-\frac{(-1)^{\alpha+n}}{(|\alpha|-n-1)!}\psi(|\alpha|-n),\quad n<|\alpha|\in{\mathbb{N}}.$ (95)
@@
$\quad\frac{1}{n!}\sum_{k=1}^{n}\mathop{\mathrm{rv}_{0}}\Bigg(\Bigg(\sum_{j=0}^{|\alpha|-n-1}+\sum_{j=|\alpha|-k}^{\infty}\Bigg)\sum_{i=0}^{n+1-k}\frac{(-1)^{j}}{j!}(|\alpha|-n-j)_{n-k}$ (96)
$\quad\times\frac{(|\alpha|-n)_{1+n-k-i}}{i!(n-i-k+1)!}z^{i+k-|\alpha|+j}\Bigg).$
@@
$G(\alpha)$ $=\operatorname{diag}\Big(\frac{\Gamma(1+\alpha+n)}{n!}\Big),$ $\quad\alpha\in{\mathbb{C}}\setminus-{\mathbb{N}},$ (98)
@@
In particular, all entries of the rows and columns $0,\dots,|\alpha|-1$ are non-zero and the sign of the entries of these rows and columns oscillates.
Integration by parts seems to be the simplest way to determine bilinear generalized integrals of Laguerre polynomials. Alternatively, we can derive Thms 3.2., 3.2. and 3.2. from Thm 2.8., using the fact that Laguerre polynomials are essentially special cases of Tricomi functions. For Thms 3.2. and 3.2. this is easy. To derive Thm 3.2. we need to analyze certain seemingly singular expressions.
Alternative proof of Thm. 3.2..
@@
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)L_{m}^{\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z$ (99a)
$=$ $\frac{(-1)^{m+n}}{m!n!}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n,\alpha}(z)U_{-1-\alpha-2m,\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z,\quad m\neq n,$ (99b)
@@
$\mathrm{gen}\int_{0}^{\infty}L_{n}^{\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z$ (100a)
$=\frac{1}{(n!)^{2}}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z.$ (100b)
The generalized integrals of Tricomi functions given by Thm 2.8. are seemingly singular for the parameters in (99b) and (100b). However, these singularities are merely apparent. To see this, we replace (99b) and (100b) by the following limits:
$I_{1}$ $\coloneqq\frac{(-1)^{m+n}}{m!n!}\lim_{\epsilon\to 0}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n-2\epsilon,\alpha}(z)U_{-1-\alpha-2m-2\epsilon,\alpha}(z){\rm e}^{-z}z^{\alpha}{\rm d}z,\quad m\neq n,$ (101)
$I_{2}$ $\coloneqq\frac{1}{(n!)^{2}}\lim_{\epsilon\to 0}\mathrm{gen}\int_{0}^{\infty}U_{-1-\alpha-2n-2\epsilon,\alpha}(z)^{2}{\rm e}^{-z}z^{\alpha}{\rm d}z.$ (102)
@@
$\times\Big(-\psi(|\alpha|-k)-\psi(k+1)+\tfrac{1}{2}H_{k}\big(\tfrac{1+\theta_{1}-|\alpha|}{2}\big)-\tfrac{1}{2}H_{|\alpha|-1-k}\big(\tfrac{3+\theta_{2}-|\alpha|}{2}+k\big)\Big)\Bigg|<\infty$ (103)
for any values of the parameters because possible vanishing denominators of the harmonic numbers are balanced by the Pochhammer symbol. Due to the $\Gamma$ functions in the prefactor, see Thm. 2.8., this sum only contributes if $\alpha>m$ and $\alpha>n$. In the cases where the sum does not contribute, we have
$I_{1}$ $=\frac{(-1)^{\alpha+m+n}}{2(m-n)n!m!}$ (104)
$\quad\times\lim_{\epsilon\to 0}\begin{cases}\dfrac{\psi(-n-\epsilon)+\psi(-\alpha-n-\epsilon)}{\Gamma(-\alpha-n-\epsilon)\Gamma(-m-\epsilon)}-(m\leftrightarrow n),&\quad\alpha\geq 0;\vskip 12.0pt\\ \dfrac{\psi(-n-\epsilon)+\psi(|\alpha|-n-\epsilon)}{\Gamma(-n-\epsilon)\Gamma(|\alpha|-m-\epsilon)}-(m\leftrightarrow n),&\quad\alpha<0,\;m\geq|\alpha|.\end{cases}$
@@
$I_{1}$ $=\begin{cases}0,&\quad\alpha\geq 0;\vskip 12.0pt\\ \dfrac{(-1)^{\alpha+n}}{(m-n)n!\Gamma(|\alpha|-n)},&\quad\alpha<0,\;m\geq|\alpha|.\end{cases}$ (106)
@@
Let us now assume that $\alpha>m$ and $\alpha>n$ (and still, $m>n$). Then
$I_{1}$ $=\frac{(-1)^{\alpha+m+n}}{2n!m!\Gamma(|\alpha|-m)\Gamma(|\alpha|-n)}\Bigg(\frac{(-1)^{m}m!\Gamma(|\alpha|-m)-(-1)^{n}n!\Gamma(|\alpha|-n)}{m-n}$ (107)
@@
$(z)_{k}H_{k}(z)=\sum_{j=0}^{k-1}\prod_{l=0,l\neq j}^{k-1}(z+l).$ (108)
With similar manipulations as in the proof of Theorem 3.2. and using Lemma B, the sum in the last line of (107) can be simplified and we obtain the expression from Thm 3.2..
The analysis of (102) is similar. We distinguish the three cases $\alpha>0$, $(\alpha<0\land n\geq|\alpha|)$ and $\alpha>n$. The sum in (54) only contributes if $\alpha>n$. To obtain the limit, one needs to evaluate certain combinations of $\psi^{\prime}(z)$, $\psi(z)$ and $\Gamma(z)$ at negative integers. We derive the respective formulas in Lemma B. ∎